OpenStudy (anonymous):

how many solutions are there to the equation 9x-16=9x-23

OpenStudy (australopithecus):

Zero solutions: 9x - 16 = 9x - 23 9x - 9x - 16 = -23 -16 =/= -23

OpenStudy (australopithecus):

This is because neither of the lines intersect

OpenStudy (australopithecus):

please see the link, if a solution existed the two lines would touch at a point

OpenStudy (australopithecus):

or many

OpenStudy (australopithecus):

there lines never touch so there is no solution, what are you doing with this is finding where the two functions or lines intersect